# Towards Quantum Accelerated Large-scale Topology Optimization

### v1

@@ 3 Proposed Methodology
In the original DVTO-MT method [20], the master problems (11) or (12) are solved by directly calling a classical MILP solver, such as Gurobi [42]. While this approach is feasible for smaller 2D designs, the MILP solver’s exponentially increasing cost can render it intractable for large-scale 3D problems, as demonstrated by our numerical tests in Sections 4.2 and 4.3.
@@ 3.1 Modified Dantzig-Wolfe (MDW) Decomposition
The mass constraint $H_{M_{0}}(\bm{\rho})\leq\bar{M}_{\text{max}}$ in Eq. (8) and the cuts $\widetilde{f}^{j}(\bm{\rho})\leq\eta$ and the trust regions $t^{j}(\bm{\rho})\leq d^{j}$ in Eq. (11) involve all design variables, and hence are categorized as global constraints. Whereas, the material usage constraint $H_{M_{e}}(\bm{\rho})\leq 1$ in (8) only applies to $n_{M}$ variables on each element, and hence can be regarded as a local constraint. As a result, the programming in Eq. (11) or (12) shows a block-angular structure, and thereby can be decomposed into a series of reduced-size sub-problems through the proposed MDW decomposition. Similar to the standard DW decomposition [33], the reduced-size sub-problems are categorized as global or local, corresponding to global and local constraints, respectively. However, we cannot simply apply the standard DW decomposition, because it typically collects all extreme points, which are generated from the local sub-problems, into the global sub-problem [33]. The “extreme points” herein refer to the points satisfying the constraints of the local sub-problems. For the programming in Eq. (11) or (12) associated with the TO problems, the number of extreme points generated from each local sub-problem can scale exponentially with respect to the number of design variables involved in the local sub-problem. As the number of design variables increases, the resultant global sub-problem could become intractable. Therefore, different from the standard DW decomposition, our proposed MDW decomposition generates extreme points iteratively by updating the local sub-problems with the Lagrangian multipliers evaluated from the global sub-problem. Instead of preparing the extreme points from the local sub-problems, we search for the optimal Lagrangian multipliers as the size of Lagrangian multipliers would not scale with the number of design variables. The Lagrangian multipliers can be well initiated from a reduced-size problem (corresponding to employing a lower discretization resolution) given the TO problem is constrained by valid physical constraints, as discussed later in §3.1.3. Given the topological similarity across different discretization resolutions, only a limited number of elements need to be modified compared to the solution obtained from a lower discretization resolution, suggesting that only a small set of extreme points are required to find the optimal solution. This MDW decomposition effectively controls the size of the global sub-problem, which in turn ensures efficiency when employed to solve the master problem (11) or (12). Although proposed in the context of TO problems, this MDW decomposition is generally applicable to large-scale binary programming with non-overlapping local constraints whose size would not scale with respect to the number of variables.
@@
$\mathcal{D}_{i}=\left\{e\left|\left\lfloor i\cdot\dfrac{n_{e}}{n_{\mathcal{D}}}\right\rfloor\leq e^{\prime}\leq\left\lfloor(i+1)\cdot\dfrac{n_{e}}{n_{\mathcal{D}}}\right\rfloor\right.\right\}\;.$ (20)
@@ 3.1.1 Local and Global Sub-problem Formulations
$\min_{\bm{\rho}_{i},\eta_{i}}\quad\eta_{i}$ $+\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}(\widetilde{f}^{j}(\bm{\rho}_{i})-\eta_{i})$ (22)
$+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$
@@
$\min_{\bm{\rho}}$ $\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}\widetilde{f}^{j}(\bm{\rho}_{i})+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$ (23)
@@
where $\bar{\bm{\rho}}=\{\bar{\rho}_{e,m}\}$ with $\bar{\rho}_{e,m}=\sum_{q=1}^{l}\rho_{e,m}^{k,l}\lambda_{i}^{q}$, $\forall e\in\mathcal{D}_{i},m=1,\dots n_{M}$, and $i=1,2,\dots,n_{\mathcal{D}}$. It is a linear programming problem, with both $\bm{\lambda}$ and $\eta$ continuous variables, and can be easily solved by a classical linear programming solver, e.g., Gurobi Optimizer [42] used in this work. During the solution of the global sub-problem (24), the Lagrange multipliers $\pi^{f}_{j}$, $\pi^{t}_{j}$, and $\pi^{M}$ are determined concurrently with $\bm{\lambda}$ and $\eta$ by the linear programming solver. These Lagrange multipliers correspond to the conversion of the three inequality constraints in (24) into equality constraints and are collectively denoted as the vector $\bm{\pi}$.
@@ 3.1.2 Initialization
To initiate the MDW decomposition, an initial guess of $\bm{\rho}$ to the programming (11) or (12) is necessary. Thus, we propose solving a reduced-size problem with a clustered design variable $\hat{\bm{\rho}}\in\{0,1\}^{n_{\mathcal{D}}\times n_{M}}$ to obtain the initial guess. Specifically, we assign all design variables $\rho_{e,m}$ in the $i$-th block with a single value of $\hat{\rho}_{i,m}$, i.e., $\rho_{e,m}=\hat{\rho}_{i,m},\quad\forall e\in\mathcal{D}_{i}$. As such, the design variables $\rho_{e,m}$ belong to the $i$-th block are aggregated into one variable $\hat{\rho}_{i,m}$, and correspondingly the coefficients associated with the design variables in (11) or (12) are also aggregated. The resultant reduced-size multi-cut problem reads:
$\min_{\hat{\bm{\rho}},\eta}$ $\eta$ (28)
s.t. $f(\bm{\rho}^{j},\mathbf{u}^{j})+\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left(\sum_{e\in\mathcal{D}_{i}}\widetilde{w}_{e,m}^{j}\right)\hat{\rho}_{i,m}$
$\quad-\sum_{e=1}^{n_{e}}\sum_{m=1}^{n_{M}}\widetilde{w}_{e,m}^{j}\rho_{e,m}^{j}\leq\eta,\quad\forall j\in\mathcal{P}_{s}$
$\dfrac{1}{n_{e}}\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left[\sum_{e\in\mathcal{D}_{i}}\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)\right]\hat{\rho}_{i,m}$
$\quad+\sum_{e=1}^{n_{e}}\left(\rho_{e,m}^{j}\right)^{2}\leq d^{j},\quad\forall j\in\mathcal{P}_{s}$
@@
$\min_{\hat{\bm{\rho}}}$ $\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left(\sum_{e\in\mathcal{D}_{i}}\widetilde{w}_{e,m}^{k-1}\right)\hat{\rho}_{i,m}$ (29)
s.t. $\dfrac{1}{n_{e}}\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left[\sum_{e\in\mathcal{D}_{i}}\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{k-1}\right)\right]\hat{\rho}_{i,m}$
$\quad+\sum_{e=1}^{n_{e}}\left(\rho_{e,m}^{k-1}\right)^{2}\leq d^{k-1}$
@@
Both can be solved by a standard MILP solver. In case that the initialization problem (28) or (29) is infeasible, $n_{\mathcal{D}}$ will be doubled, and (28) or (29) will then be reformulated with the new $n_{\mathcal{D}}$. This process can be repeated until a feasible solution to the initialization problem is found. Our numerical experiments indicate that infeasible solutions are uncommon during the initialization stage, provided $n_{\mathcal{D}}$ is not set too small. For all numerical experiments in this work, we set $n_{\mathcal{D}}=100$. The solution to the initialization problem yields the initial Lagrange multiplier values, which are then used when solving the local sub-problems in the first iteration step.
@@ 3.1.3 Iteration Process and Algorithm
Here, the vector $\bm{\pi}$ is defined as: $\bm{\pi}=\{\pi^{f},\pi^{t},\pi^{M}\}$ if the original programming is multi-cut as in (11), or $\bm{\pi}=\{\pi^{t},\pi^{M}\}$ if the original programming is single-cut as in (12).
@@ A QUBO problem generally reads:
[…]
@@
$\sum_{e\in\mathcal{D}_{i}}\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}\widetilde{f}^{j}(\bm{\rho}_{i})+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$ (33)
$+$ $\Pi\sum_{e\in\mathcal{D}_{i}}\left[\sum_{m=1}^{n_{M}}\rho_{e,m}+\phi_{e}-1\right]^{2}\;,$
@@
$\sum_{e\in\mathcal{D}_{i}}\sum_{m=1}^{n_{M}}\widetilde{w}_{e,m}^{k-1}\rho_{e,m}+\pi^{t}t^{k-1}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$ (34)
$+$ $\Pi\sum_{e\in\mathcal{D}_{i}}\left[\sum_{m=1}^{n_{M}}\rho_{e,m}+\phi_{e}-1\right]^{2}\;.$
@@ 4.1 Problem Setup
The design domain is a 3D box, with the length of $L=40$ m, the width of $W=10$ m, and the height of $H=10$ m, as depicted in Figure 1. A solid plate with the thickness of 0.4 m is placed on top of the box, designated by the red region in Figure 1. This top plate region is excluded from the TO design, and hence is always assigned solid elements of the strongest material. The design domain is fixed in two striped regions at either end of the bottom boundary, each with a width of 2 m, as designated by the orange regions in Figure 1. The gray region represents the actual design domain, allowing for free-form topologies and the use of any candidate materials, as determined by the optimization solution. A uniformly distributed loading force with the magnitude of $-10^{3}N/m^{2}$ is exerted on the top boundary of the design domain, as indicated by blue arrows in Figure 1. Three different discretization resolutions are examined, including $50\times 200\times 50$, $100\times 400\times 100$, and $150\times 600\times 150$.
@@ 4.2 Single-Material Design
Resolution MDW-CC SIMP ($N_{\text{Iter}}=200$) SIMP ($N_{\text{Iter}}=400$)
@@
As discussed in §3.2, the local sub-problems (22) or (26) can be solved using quantum computing to leverage its advantages in solving BIP problems. Thus, in the second-tier validation, we verify the QUBO formulated for the local sub-problems. Due to the practical limitations of accessing real quantum hardware, we opted to solve the QUBO in (33) with the QAOA algorithm, leveraging the simulator provided by cuda-quantum [45]. Given the limited number of qubits that can be simulated on a classical computer, we constrained the size of $|\mathcal{D}_{i}|$ up to 8 for validation purposes. Thirty local sub-problems were randomly selected from the case with the discretization resolution of $50\times 200\times 50$. The number of layers in the ansatz was set to $\mathcal{L}=4$ or $\mathcal{L}=5$. To evaluate accuracy, we compared the QUBO’s solution generated by QAOA against that yield by calling the classical optimizer [42] for solving the original local sub-problem. The accuracy is evaluated by the percentage of agreement between the two solutions across all 30 problems solved, as presented in Table 4. The results show that with a sufficient number of layers in the ansatz, specifically $\mathcal{L}=5$, our QUBO formulation successfully guides the optimization to the desired solution, thus validating its use for updating design variables in the TO process.

### v2

@@ 3 Proposed Methodology
In the original DVTO-MT framework [20], the master problems (11) and (12) are solved by directly calling a classical MILP solver from Gurobi [44], which relies on a branch-and-bound approach. While this approach is feasible for smaller 2D designs, the MILP solver’s exponentially increasing cost can render it intractable for large-scale 3D problems, as demonstrated by our numerical tests in Sections 4.2 and 4.3.
@@ 3.1 Modified Dantzig-Wolfe (MDW) Decomposition
The mass constraint $H_{M_{0}}(\boldsymbol{\rho})\leq\bar{M}_{\text{max}}$ in Eq. (8) and the cuts $\widetilde{f}^{j}(\bm{\rho})\leq\eta$ and the trust regions $t^{j}(\bm{\rho})\leq d^{j}$ in Eq. (11) involve all design variables, and hence are categorized as global constraints. Whereas, the material usage constraint $H_{M_{e}}(\bm{\rho})\leq 1$ in (8) only applies to $n_{M}$ variables on each element, and hence can be regarded as a local constraint. As a result, the programming in Eq. (11) or (12) shows a block-angular structure, and thereby can be decomposed into a series of reduced-size sub-problems through the proposed MDW decomposition. Similar to the standard DW decomposition [33], the reduced-size sub-problems are categorized as global or local, corresponding to global and local constraints, respectively. However, we cannot simply apply the standard DW decomposition, because it typically collects all extreme points, which are generated from the local sub-problems, into the global sub-problem [33]. The “extreme points” herein refer to the points satisfying the constraints of the local sub-problems. For the programming in Eq. (11) or (12) associated with the TO problems, the number of extreme points generated from each local sub-problem scales exponentially with respect to the number of design variables involved in the local sub-problem. As the number of design variables increases, the resultant global sub-problem could become intractable. Therefore, different from the standard DW decomposition, our proposed MDW decomposition generates extreme points iteratively by updating the local sub-problems with the Lagrangian multipliers evaluated from the global sub-problem. Instead of preparing the extreme points from the local sub-problems, we search for the optimal Lagrangian multipliers as the size of Lagrangian multipliers would not scale with the number of design variables. The Lagrangian multipliers can be well initiated from a reduced-size problem (corresponding to employing a lower discretization resolution) given the TO problem is constrained by valid physical constraints, as discussed later in §3.1.3. Given the topological similarity across different discretization resolutions, only a limited number of elements need to be modified compared to the solution obtained from a lower discretization resolution, suggesting that only a small set of extreme points are required to find the optimal solution. This MDW decomposition effectively controls the size of the global sub-problem, which in turn ensures efficiency when employed to solve the master problem (11) or (12). Although proposed in the context of TO problems, this MDW decomposition is generally applicable to large-scale binary programming with non-overlapping local constraints whose size would not scale with respect to the number of variables.
@@
Eq. (20) implies that the $i$-th block collects the indices of the elements $e$ with the same or close value of $\widetilde{\rho}_{e}$, where $\lfloor\cdot\rfloor$ denoting the floor function. As such, each block manages about $\dfrac{n_{e}n_{M}}{n_{\mathcal{D}}}$ design variables. The vector $\bm{\rho}_{i}$ is formed such that $\bm{\rho}_{i}=\left\{\rho_{e,m}|\forall e\in\mathcal{D}_{i},m=1,\dots n_{M}\right\}$. This indexing and partitioning strategy can ensure that all design variables $\rho_{e,m}$ associated with a given element $e$ are grouped into the same block and thus included within the same local sub-problem. This is critical for multi-material design problems, as partitioning the design variables defined on the same element into different blocks would no longer preserve the block-angular structure of the problem which is required by the proposed MDW decomposition method. Alternative partitioning strategies, such as spatially-based partitioning, may also be employed, provided they guarantee that all design variables associated with the same element remain within the same block. For single-material design problems, however, only one design variable is defined per element, so the block-angular structure is inherently preserved regardless of the partitioning strategy.
@@ 3.1.1 Local and Global Sub-problem Formulations
$\min_{\boldsymbol{\rho}_{i},\eta_{i}}$ $\eta_{i}+\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}(\widetilde{f}^{j}(\bm{\rho}_{i})-\eta_{i})+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$ (22)
@@
$\min_{\boldsymbol{\rho}}$ $\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}\widetilde{f}^{j}(\bm{\rho}_{i})+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})$ (23)
@@
where $\bar{\bm{\rho}}=\{\bar{\rho}_{e,m}\}$ with $\bar{\rho}_{e,m}=\sum_{q=1}^{l}\rho_{e,m}^{k,l}\lambda_{i}^{q}$, $\forall e\in\mathcal{D}_{i},m=1,\dots n_{M}$, and $i=1,2,\dots,n_{\mathcal{D}}$. It is a linear programming problem, with both $\bm{\lambda}$ and $\eta$ continuous variables, and can be easily solved by a classical linear programming solver, e.g., Gurobi Optimizer [44] used in this work. During the solution of the global sub-problem (24), the Lagrange multipliers $\pi^{f}_{j}$, $\pi^{t}_{j}$, and $\pi^{M}$ are determined concurrently with $\boldsymbol{\lambda}$ and $\eta$ by the linear programming solver. These Lagrange multipliers correspond to the conversion of the three inequality constraints in (24) into equality constraints and are collectively denoted as the vector $\bm{\pi}$.
@@ 3.1.2 Initialization
To initiate the MDW decomposition, an initial guess of $\boldsymbol{\rho}$ to the programming (11) or (12) is necessary. Thus, we propose solving a reduced-size problem with a clustered design variable $\hat{\bm{\rho}}\in\{0,1\}^{n_{\mathcal{D}}\times n_{M}}$ to obtain the initial guess. Specifically, we assign all design variables $\rho_{e,m}$ in the $i$-th block with a single value of $\hat{\rho}_{i,m}$, i.e., $\rho_{e,m}=\hat{\rho}_{i,m},~\forall e\in\mathcal{D}_{i}$. As such, the design variables $\rho_{e,m}$ belong to the $i$-th block are aggregated into one variable $\hat{\rho}_{i,m}$, and correspondingly the coefficients associated with the design variables in (11) or (12) are also aggregated. The resultant reduced-size multi-cut problem reads:
$\min_{\hat{\boldsymbol{\rho}},\eta}$ $\eta$ (28)
s.t. $f(\bm{\rho}^{j},\mathbf{u}^{j})+\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left(\sum_{e\in\mathcal{D}_{i}}\widetilde{w}_{e,m}^{j}\right)\hat{\rho}_{i,m}-\sum_{e=1}^{n_{e}}\sum_{m=1}^{n_{M}}\widetilde{w}_{e,m}^{j}\rho_{e,m}^{j}\leq\eta,\quad\forall j\in\mathcal{P}_{s}$
$\dfrac{1}{n_{e}}\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left[\sum_{e\in\mathcal{D}_{i}}\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{j}\right)\right]\hat{\rho}_{i,m}+\sum_{e=1}^{n_{e}}\left(\rho_{e,m}^{j}\right)^{2}\leq d^{j},\quad\forall j\in\mathcal{P}_{s}$
@@
$\min_{\hat{\boldsymbol{\rho}}}$ $\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left(\sum_{e\in\mathcal{D}_{i}}\widetilde{w}_{e,m}^{k-1}\right)\hat{\rho}_{i,m}$ (29)
s.t. $\dfrac{1}{n_{e}}\sum_{m=1}^{n_{M}}\sum_{i=1}^{n_{\mathcal{D}}}\left[\sum_{e\in\mathcal{D}_{i}}\left(1-2\sum_{m=1}^{n_{M}}\rho_{e,m}^{k-1}\right)\right]\hat{\rho}_{i,m}+\sum_{e=1}^{n_{e}}\left(\rho_{e,m}^{k-1}\right)^{2}\leq d^{k-1}$
@@
Both can be solved by a standard MILP solver. In case that the initialization problem (28) or (29) is infeasible, $n_{\mathcal{D}}$ will be doubled, (28) or (29) will then be reformulated with the new $n_{\mathcal{D}}$. This process can be repeated until a feasible solution to the initialization problem is found. The rationale behind this process is that the reduced-sized problem imposes a restriction to the original problem as it constrains all design variables within a block to take the same value. Doubling the number of blocks, or equivalently halving the number of design variables within each block, thus relaxes this constraint and in turn increases the likelihood that the resulting problem is feasible. Our numerical experiments indicate that infeasible solutions are uncommon during the initialization stage, provided $n_{\mathcal{D}}$ is not set too small. For all numerical experiments in this work, we set $n_{\mathcal{D}}=100$. The solution to the initialization problem yields the initial Lagrange multiplier values, which are then used when solving the local sub-problems in the first iteration step.
@@ 3.1.3 Iteration Process and Algorithm
Here, the vector $\bm{\pi}$ is defined as: $\bm{\pi}=\{\pi^{f},\pi^{t},\pi^{M}\}$ if the original programming is multi-cut as in (11), or $\bm{\pi}=\{\pi^{t},\pi^{M}\}$ if the original programming is single-cut as in (12). The tolerance in Eq. (30) is consistently set to $10^{-6}$ throughout all numerical results presented in the main text. The robustness of the proposed method with respect to this parameter is discussed in A.
@@ A QUBO problem generally reads:
$\begin{aligned} &\mathbf{\mathbb{Z}}^{*}=\min_{\mathbf{{\mathbb{Z}}}\in\{0,1\}^{n}}\mathcal{J}\\ &\mathcal{J}=\mathbf{{\mathbb{Z}}}^{\intercal}\mathbf{Q}\mathbf{{\mathbb{Z}}}=\sum_{i=1}^{n}Q_{ii}\mathbb{Z}_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}Q_{ij}\mathbb{Z}_{i}\mathbb{Z}_{j}\end{aligned}\;,$ (31)
@@
$\sum_{e\in\mathcal{D}_{i}}\sum_{j\in\mathcal{P}_{s}}\pi^{f}_{j}\widetilde{f}^{j}(\bm{\rho}_{i})+\sum_{j\in\mathcal{P}_{s}}\pi_{j}^{t}t^{j}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})+\Pi\sum_{e\in\mathcal{D}_{i}}\left[\sum_{m=1}^{n_{M}}\rho_{e,m}+\phi_{e}-1\right]^{2}\;,$ (33)
@@
$\sum_{e\in\mathcal{D}_{i}}\sum_{m=1}^{n_{M}}\widetilde{w}_{e,m}^{k-1}\rho_{e,m}+\pi^{t}t^{k-1}(\bm{\rho}_{i})+\pi^{M}H_{M_{0}}(\bm{\rho}_{i})+\Pi\sum_{e\in\mathcal{D}_{i}}\left[\sum_{m=1}^{n_{M}}\rho_{e,m}+\phi_{e}-1\right]^{2}\;.$ (34)
@@ 4.1 Problem Setup
The design domain is a 3D box, with the length of $L=40$ m, the width of $W=10$ m, and the height of $H=10$ m, as depicted in Figure 1. A solid plate with the thickness of 0.4 m is placed on top of the box, designated by the red region in Figure 1. This top plate region is excluded from the TO design, and hence is always assigned solid elements of the strongest material. The design domain is fixed in two striped regions at either end of the bottom boundary, each with a width of 2 m, as designated by the orange regions in Figure 1. The gray region represents the actual design domain, allowing for free-form topologies and the use of any candidate materials, as determined by the optimization solution. A uniformly distributed loading force with the magnitude of $10^{3}N/m^{2}$ is exerted on the top boundary of the design domain, as indicated by blue arrows in Figure 1. Three different discretization resolutions are examined, including $50\times 200\times 50$, $100\times 400\times 100$, and $150\times 600\times 150$.
@@ 4.2 Single-Material Design
Resolution MDW-$C$ SIMP ($N_{\text{Iter}}=200$) SIMP ($N_{\text{Iter}}=400$)
@@
[…]
$\mathbb{Z}_{i}=\dfrac{\zeta_{i}+1}{2},\quad\mathbb{Z}_{i}\in\{0,1\},\quad\zeta_{i}\in\{-1,1\}\;.$ (37)
The objective function $\mathcal{J}$ can then be rewritten as:
$\mathcal{J}=\sum_{i=1}^{n}Q_{ii}\left(\dfrac{\zeta_{i}+1}{2}\right)+\sum_{i=1}^{n}\sum_{j=i+1}^{n}Q_{ij}\left(\dfrac{\zeta_{i}+1}{2}\right)\left(\dfrac{\zeta_{j}+1}{2}\right)~~\to~~\mathcal{J}^{\prime}=\sum_{i=1}^{n}\dfrac{1}{2}\left({Q_{ii}+\sum_{j=i+1}^{n}Q_{ij}}\right)\zeta_{i}+\sum_{i=1}^{n}\sum_{j=i+1}^{n}\frac{1}{4}Q_{ij}\zeta_{i}\zeta_{j}$ (38)
where $\mathcal{J}^{\prime}$ omits the constants in $\mathcal{J}$. Thus, we obtained the Hamiltonian as:
[…]
Given the limited number of qubits that can be simulated on a classical computer, we constrained the size of $|\mathcal{D}_{i}|$ up to 8 for validation purposes. Thirty local sub-problems were randomly selected from the case with the discretization resolution of $50\times 200\times 50$. The number of layers in the ansatz was set to $\mathcal{L}=4$ or $\mathcal{L}=5$. To evaluate accuracy, we compared the QUBO’s solution generated by QAOA against that yield by calling the classical optimizer [44] for solving the original local sub-problem. The accuracy is evaluated by the percentage of agreement between the two solutions across all 30 problems solved, as presented in Table 4. The results show that with a sufficient number of layers in the ansatz, specifically $\mathcal{L}=5$, our QUBO formulation successfully guides the optimization to the desired solution, thus validating its use for updating design variables in the TO process.